# October 2024

### v1

@@
{mdframed}
[linecolor=magenta] [Uncaptioned image] This version is a preprint of a paper authored by K. Okamura, titled ‘Evolving interdisciplinary contributions to global societal challenges: A 50-year overview’, posted on arXiv. The supplementary materials accompanying this manuscript have been seamlessly integrated into this single file.
@@ 1 Introduction: Interdisciplinary efforts for global societal challenges
Nevertheless, there remains a lack of quantitative evidence on how interdisciplinary efforts have practically contributed to addressing global societal challenges. Several studies employing bibliometric approaches have analysed the relationship between societal challenges and various scientific subfields or relevant scientific journals, or have compared national contributions (Bedard-Vallee et al.,, 2023; Garai et al.,, 2023; Mishra et al.,, 2024; Raman et al.,, 2024; Sweileh,, 2020; Yamaguchi et al.,, 2023; Yumnam et al.,, 2024). These studies offer valuable insights into the connections between academia, the international community, and the SDGs. However, examining long-term contributions to global societal challenges from an interdisciplinary or cross-disciplinary perspective remains challenging, especially across the period encompassing both pre- and post-UN initiatives, including the MDGs and SDGs.
@@
The knowledge and strategies that are effective in addressing specific societal challenges can differ significantly depending on the nature of the issue concerned. It is entirely possible for individuals from completely different disciplines to work on the same societal challenge without recognising each other’s efforts, applying knowledge and methods unique to their respective disciplines. For instance, which research fields have contributed to SDG 5, which aims at achieving gender equality and empowering all women and girls? Fields like sociology, political science, or psychology may immediately come to mind. However, as shown later (see Figure 3), medicine has made a significant contribution, especially in recent years—a fact that may not be readily apparent compared to the aforementioned fields. Each societal challenge often involves significant contributions from disciplines that may not initially be given much consideration. Understanding this reality and its temporal trends offers a vital perspective on interdisciplinary efforts and their contributions to various societal challenges. This understanding helps reveal the organic collaborative relationships and structures between academia and the policy sphere that work towards addressing common global problems. Such insights are invaluable for shaping future collaboration across disciplines, sectors, and borders.
@@
In practice, conducting such analyses has proven challenging, primarily due to the difficulty of assessing whether—and to what extent—a given contribution has helped address, let alone resolve, a societal challenge (D’Este and Robinson-García,, 2023; Sandes-Guimarães et al.,, 2022). These evaluations are inherently part of a complex social and political process, and the methods, timeframes, and social consensus required make it difficult to discuss the results with any degree of certainty—scientifically, politically, or practically. Nevertheless, in today’s era of rapidly advancing Open Science and Data Science (Dong et al.,, 2017; Wittenburg,, 2021), and with the increasing availability of diverse scientometric data, it is now more feasible than ever to gain indirect insights through extensive data analysis. This is the approach we adopt in this paper.
@@ 2.1 The OpenAlex data
The number of publications examined in this study was $N=100,700$, derived from 1,000 publications per field per publication year, multiplied by 19 fields ($\mathscr{D}$) and 53 publication years (1970–2022). Refer to Supplementary Materials, including Tables S1 and S2, for further details on the OpenAlex data used in this study.
@@ 2.2 The field-affinity score
To calculate the distance between fields A and B in $\mathscr{D}$, denoted as $d_{\mathrm{A}\mathrm{B}}$, we first identified and counted the number of publications ($n_{\mathrm{A}\cup\mathrm{B}}$) with a positive field-affinity score for either or both fields A and B from the total of $N=100,700$ publications. We then counted the number of publications ($n_{\mathrm{A}\cap\mathrm{B}}$) that had positive field-affinity scores for both fields A and B. Finally, we calculated the ratio of these counts and subtracted it from 1 to obtain the Jaccard distance between fields A and B:
@@
Here, we quantified interdisciplinarity using the so-called effective number of disciplines (Jost,, 2006; Leinster and Cobbold,, 2012; Okamura,, 2020). This index offers advantages over other commonly used diversity measures, such as the simple ‘variety’ index, Simpson’s index, or popular entropy measures like Shannon’s, in that it not only accounts for the number of fields and their composition (i.e., relative abundances) but also appropriately considers the distances between the constituent fields. Moreover, it is the most universal, as it can mathematically accommodate both broad categorisation and finer subdivisions of constituents in a consistent manner, satisfying the so-called Nesting Principle (Okamura,, 2020), which is not upheld by other measures such as the Rao-Stirling index (Rao,, 1982).
@@ 2.3 The SDG-affinity score
[…]
@@ 4 Conclusions: Towards evidence-based actions for sustainable development
This paper has provided an overview of how various disciplines are connected to different global societal challenges and how these connections have evolved over more than half a century—an area that has previously lacked substantial quantitative evidence. Specifically, we utilised two types of scores based on specific machine learning algorithms employed in the OpenAlex platform: the field-affinity scores and the SDG-affinity scores. These scores were used to develop appropriate measures for analysing the organic and dynamic relationships between different disciplines and the SDGs.
Our analysis suggests that since the 2000s, when the MDGs were implemented in earnest, there has been a clear trend towards an increase in publications and/or citations with interdisciplinary content, with many fields showing a tendency to become more interconnected. The focus on specific global societal issues, as represented by the SDGs, has also highlighted that the interdisciplinarity of these issues and the ways in which they have changed over time are markedly different, reflecting the unique characteristics of each issue. A considerable body of literature has concentrated, and will continue to concentrate, on the analysis and discussion of each of the SDGs, particularly as we approach the five-year mark before the SDG target deadline. Within these ongoing discussions, the findings and visualisations presented in this study offer valuable and insightful perspectives from an interdisciplinary and cross-sectoral standpoint. These insights are especially relevant to policymakers and practitioners as they reflect on the past decades and strategise for the future.
Finally, we would like to address some considerations regarding the methodology employed in this study and outline expectations for future research. It is essential to continue evaluating and refining the approach used in this study, which applies an Open Bibliometrics-based lens to observe the complex landscape of diverse global societal challenges. Open data catalogues of scholarly publication data, including OpenAlex, utilised in this paper, are continually improving in both the quality and quantity of their collections, and these resources will increasingly serve as valuable tools for analysing and visualising global trends. While commercial databases like Web of Science and Scopus are beneficial for conducting precise analyses from specific perspectives, such open and freely accessible resources (Delgado-Quirós and Ortega,, 2024) can also offer valuable insights into broad trends, particularly for policy-related purposes (Kitajima and Okamura,, 2023; OECD.AI, n.d., ; Okamura,, 2023, 2024; UNESCO,, 2023). In these contexts, the emphasis is on quick comprehension and a general, albeit rough, understanding. By recognising the unique characteristics—including both strengths and limitations—of each scholarly database and by using or combining them appropriately for different applications, we can gain new insights and develop more effective lenses through which to examine global trends (Culbert et al.,, 2024; Haupka et al.,, 2024; Liu et al.,, 2023; Visser et al.,, 2021). Doing so will further enhance the quality and dynamism of discussions surrounding societal impact in both policy and academic spheres, contributing meaningfully to these critical dialogues.
@@ Acknowledgements.
[…]
@@
[…]
@@
[…]

### v2

# October 2025 [v2]

@@
[Uncaptioned image] This version reproduces the content of the paper authored by K. Okamura, titled ‘Evolving interdisciplinary contributions to global societal challenges: A 50-year overview’, published in World Development Perspectives (doi: 10.1016/j.wdp.2025.100728). The supplementary materials accompanying this manuscript have been seamlessly integrated into this single file.
@@ 1 Introduction: Interdisciplinary efforts for global societal challenges
Nevertheless, there remains a lack of systematic and long-term quantitative evidence on how interdisciplinary efforts have contributed to addressing global societal challenges. Prior bibliometric studies have mapped linkages between societal challenges and scientific subfields, profiled journal-level participation, and compared national contributions to the SDGs (Bedard-Vallee et al.,, 2023; Garai et al.,, 2023; Mishra et al.,, 2024; Raman et al.,, 2024; Sweileh,, 2020; Yamaguchi et al.,, 2023; Yumnam et al.,, 2024). Collectively, this literature delineates where disciplinary engagement is concentrated versus diffuse and how SDG-relevant outputs are distributed across fields and venues. However, most analyses focus on specific domains or limited time windows, providing only partial evidence on long-run, cross-SDG interdisciplinarity.
@@
The knowledge and strategies effective in addressing specific societal challenges can differ markedly depending on the nature of the issue. It is entirely possible for individuals—researchers and policymakers alike—from completely different disciplines and policy areas to work on the same challenge without recognising each other’s contributions, applying knowledge and methods unique to their respective fields. For instance, in the case of SDG 5, which aims to achieve gender equality and empower all women and girls, disciplines such as sociology, political science, or psychology may immediately come to mind. However, as shown later (Figure 3), medicine has also made a substantial contribution in recent years, a fact not readily apparent when compared with the aforementioned fields.
Each societal challenge therefore often involves significant contributions from disciplines that might not initially be expected. Recognising these contributions and their temporal dynamics provides a vital perspective on how interdisciplinary efforts evolve and how they contribute to addressing global challenges. Such an understanding also helps reveal the organic collaborative relationships and structures between academia and the policy sphere that work towards common global problems. These insights are invaluable for shaping future collaboration across disciplines, sectors and borders.
@@
In practice, conducting such analyses has proven challenging, primarily due to the difficulty of assessing whether, and to what extent, a given contribution has helped address, let alone resolve, a societal challenge (D’Este and Robinson-García,, 2023; Sandes-Guimarães et al.,, 2022). These evaluations are inherently part of a complex social and political process, and the methods, timeframes, and social consensus required make it hard to discuss the results with any degree of certainty—scientifically, politically or practically. Nevertheless, in today’s era of rapidly advancing Open Science and Data Science (Dong et al.,, 2017; Wittenburg,, 2021), and with the increasing availability of diverse scientometric data, it has become more feasible than ever to gain indirect insights through extensive data analysis. This is the approach we adopt in this paper.
While previous studies provide valuable insights into the relationships among scientific subfields, journals, and national or international contributions to the SDGs, important gaps remain. Most notably, we still lack systematic, longitudinal evidence on how disciplinary contributions to global societal challenges have evolved over time and the extent to which different challenges follow distinct trajectories of interdisciplinarity across the pre-MDG and post-SDG periods. Against this backdrop, the present study addresses the following guiding research question: How have the connections between research disciplines and global societal challenges, as represented by the SDGs, developed over time?
To answer this question, the paper proceeds as follows. Section 2 describes the data and methodology, focusing on two machine learning–based measures derived from the OpenAlex platform: the field-affinity scores and the SDG-affinity scores. Section 3 presents the main results, highlighting both overall trends and SDG-specific dynamics. Section 4 discusses the implications of these findings and concludes with a synthesis of the study’s contributions, limitations, and directions for future research.
@@ 2.1 The OpenAlex data
In total, the dataset comprised N=100,700N=100{,}700 individual publications. This number results from sampling 1,000 publications per field per publication year, across 19 fields ($\mathscr{D}$) and 53 publication years (1970–2022). Refer to Supplementary Materials, including Tables S1 and S2, for further details on the OpenAlex data used in this study.
@@ 2.2 The field-affinity score
To calculate the distance between fields A and B in $\mathscr{D}$, denoted as $d_{\mathrm{A}\mathrm{B}}$, we first identified and counted the number of publications ($n_{\mathrm{A}\cup\mathrm{B}}$) with a positive field-affinity score for either or both fields A and B from the total of N=100,700N=100{,}700 publications. We then counted the number of publications ($n_{\mathrm{A}\cap\mathrm{B}}$) that had positive field-affinity scores for both fields A and B. Finally, we calculated the ratio of these counts and subtracted it from 1 to obtain the Jaccard distance between fields A and B:
@@
Here, we quantified interdisciplinarity using the so-called effective number of disciplines (Jost,, 2006; Leinster and Cobbold,, 2012; Okamura,, 2020). This index offers advantages over other commonly used diversity measures, such as the simple ‘variety’ index, Simpson’s index, or popular entropy measures like Shannon’s, in that it not only accounts for the number of fields and their composition (i.e. relative abundances) but also appropriately considers the distances between the constituent fields. Moreover, it is the most universal, as it can mathematically accommodate both broad categorisation and finer subdivisions of constituents in a consistent manner, satisfying the so-called Nesting Principle (Okamura,, 2020), which is not upheld by other measures such as the Rao–Stirling index (Rao,, 1982).
@@ 2.3 The SDG-affinity score
In addition to the field-affinity scores, our compiled OpenAlex data also contained a score indicating the affinity of each publication to each of the 17 SDGs, which we refer to as the SDG-affinity score. [^3] https://docs.openalex.org/api-entities/works/work-object#sustainable_development_goals. The scores are calculated using specific machine learning algorithms (Jaworeck,, 2022; Vanderfeesten et al.,, 2021). This score is assigned to each publication for each SDG, labelled by $m=1,\,\dots,\,17\,(\equiv M)$, as a number between 0 and 1. A score of 0 indicates that a publication is not related to the SDG, while a score of 1 indicates maximal relevance to the SDG.
@@ 4 Conclusions: Towards evidence-based actions for sustainable development
This paper has provided an overview of how various disciplines are connected to different global societal challenges and how these connections have evolved over more than half a century, an area that has previously lacked substantial quantitative evidence. Specifically, we employed two types of scores based on specific machine learning algorithms in the OpenAlex platform: field-affinity scores and SDG-affinity scores. These measures allowed us to analyse the organic and dynamic relationships between different disciplines and the SDGs.
Our research question asked how the connections between research disciplines and global societal challenges—here represented by the SDGs—have developed over time. Our analysis provides a clear answer: since the 2000s, coinciding with the implementation of the MDGs, interdisciplinary publications and citations have increased markedly, with many fields showing a tendency to become more interconnected. At the same time, the evolution of interdisciplinarity has varied significantly across SDGs, reflecting the unique characteristics of each challenge. In this sense, the study contributes novel quantitative evidence of both the overall strengthening of interdisciplinarity and the heterogeneous dynamics shaping specific global issues.
Our analysis also reveals striking asymmetries across specific SDGs. For example, SDG3 has consistently been dominated by Medicine, whereas other goals reflect more balanced contributions from a wide range of disciplines. Moreover, the degree of interdisciplinarity itself shows diverse trajectories, with some SDGs exhibit increasing integration over time, others show little change, and a few even display declining trends, reflecting the distinctive characteristics of each societal challenge. These extreme results underscore the heterogeneous and dynamic nature of interdisciplinarity: while certain global issues are addressed primarily through the lens of a single discipline, others require, and indeed demonstrate, a more diversified set of disciplinary engagements, with compositions that continue to evolve over time.
Beyond its scholarly contribution, the study also offers important practical implications. Tackling complex societal challenges requires not only the integration of the latest scientific knowledge and emerging technological trends, such as the rapid incorporation of AI, but also effective responses to major initiatives spearheaded by influential countries and international organisations. These findings underscore the need for funding agencies and policymakers, working closely with international research communities, to sustain and expand efforts that foster interdisciplinarity. Such efforts should be internationally coordinated and supported by continually updated technological infrastructures, ensuring that the growing interdisciplinarity observed in academic research translates into more effective real-world solutions.
In particular, funding agencies and policymakers may consider three complementary measures: (i) establishing dedicated funding programmes that explicitly target interdisciplinary research related to the SDGs—and their successors in the post-2030 agenda—in the era of advanced AI platforms; (ii) issuing calls for proposals that require or strongly encourage cooperation across fields, including the Arts, Humanities, and Social Sciences, where appropriate; and (iii) incentivising cross-border partnerships between academia, industry, and policy sectors through effective and future-oriented science diplomacy where needed. Taken together, these measures would help ensure that the growing interdisciplinarity observed in academic research is effectively channelled into real-world solutions to global societal challenges.
As we approach the five-year mark before the SDG target deadline, the insights and visualisations presented here provide valuable perspectives from an interdisciplinary and cross-sectoral standpoint. They offer both researchers and decision-makers a clearer understanding of how disciplinary boundaries have shifted in response to global challenges, while also highlighting concrete pathways for advancing collaboration and evidence-based action in the years to come.
Finally, this study highlights several methodological considerations and limitations that also point to directions for future research. Because the analysis relies on OpenAlex, the findings inevitably reflect the characteristics of this database. Although coverage is expanding, it remains uneven across disciplines, regions, and languages: English-language outputs are generally overrepresented, while publications in local languages or regional journals are less visible. Nevertheless, compared with other major databases, OpenAlex offers relatively better representation of non-English publications and has advantages in several other respects (Priem et al.,, 2022; Delgado-Quirós and Ortega,, 2024). In particular, it provides complementary strengths to commercial databases, especially for policy-related perspectives where quick and broad overviews are essential (Kitajima and Okamura,, 2025; OECD.AI, n.d., ; Okamura,, 2023, 2024; UNESCO,, 2023). Recognising the unique features of each database, including their strengths and limitations, and combining them appropriately can generate richer insights into global research dynamics (Culbert et al.,, 2025; Haupka et al.,, 2024; Liu et al.,, 2023; Visser et al.,, 2021).
A further limitation concerns the field-classification scheme employed, since alternative specifications could yield different distributions of interdisciplinarity and collaboration. Acknowledging these constraints clarifies the scope of the present contribution and highlights the importance of integrating complementary data sources and refining classification methods in future work. In this way, bibliometric analyses can continue to evolve as a crucial tool for examining the interplay between research and global societal challenges, thereby strengthening both scholarly understanding and policy-oriented debates. By fostering such methodological refinement, future work can further enhance the quality and dynamism of discussions on societal impact in both policy and academic spheres, contributing meaningfully to these critical dialogues.
@@ Acknowledgements.
Table S1: The data contained in the ‘Suppl_data_1.csv’ file. Domains are labelled by $\texttt{{\color[rgb]{1,0,0}$\alpha$}}=\texttt{1},\,\dots,\,\texttt{4}$, where 1: Life Sciences, 2: Social Sciences, 3: Physical Sciences, and 4: Health Sciences.
@@
Table S2: The data contained in the ‘Suppl_data_2.csv’ file. Disciplines are labelled by $\text{{\color[rgb]{1,0,0}k}}=1,\,\dots,\,19$, where 1: Political Science, 2: Philosophy, 3: Economics, 4: Business, 5: Psychology, 6: Mathematics, 7: Medicine, 8: Biology, 9: Computer Science, 10: Geology, 11: Chemistry, 12: Art, 13: Sociology, 14: Engineering, 15: Geography, 16: History, 17: Materials Science, 18: Physics, and 19: Environmental Science.
@@
discipk Score of affinity to each of the level-0 concepts defined in OpenAlex ($\text{{\color[rgb]{1,0,0}k}}=1,\,\dots,\,19$)